What is the least common denominator for these four fractions,3\4,4\5,1\2,and 1\10? Please show me the work.
Answer by stanbon(75887)

What is the least common denominator for these four fractions,
3\4,4\5,1\2,and 1\10? Please show me the work.
===============
Prime-factor form of the denominators:
4 = 2^2
5 = 5
2 = 2
10 = 2*5
===========
lcm = 2^2*5 20
===============
Cheers,
Stan H.
==========